OpenStudy (anonymous):

In a triangle, the second angle is twice the measure of the first, and the third angle is three times the measure of the second. Find the measure of the third angle. A. 20 degrees B. 30 degrees C. 120 degrees D. 180 degrees

OpenStudy (lacypennelll):

http://www.algebra.com/algebra/homework/Angles/Angles.faq.question.227900.html try this hun ^-^

OpenStudy (zehanz):

Suppose the value of the first angle is called a. The second is twice the first, so it's 2a. The third is three times the second, so it's .... Adding all three of them up gives: xxx times a = 180. You now know a, so you can calculate the measure of the third angle!

OpenStudy (anonymous):

the equation would 2a + 3a = 180

OpenStudy (zehanz):

What about the first angle with measure a?

OpenStudy (anonymous):

6

OpenStudy (zehanz):

It is a + 2a + 3*(2a) = 180

OpenStudy (anonymous):

oh. so the rest would be 2a + 6a = 180

OpenStudy (zehanz):

In total: 9a = 180, so a = 20, so the measure of the third is...

OpenStudy (anonymous):

20

OpenStudy (zehanz):

No, that is a, the measure of the first one! The third one is 6a = 6*20 = 120, so it is answer C.